Find the y- intercept
-2x - y = -10

Answer by nerdybill(7384)

Find the y- intercept
-2x - y = -10
.
Put the above equation into the "slope-intercept" form:
y = mx + b
where
m is the slope
b is the y-intercept at (0,b)
.
-2x = y -10
-2x + 10 = y
So,
m is -2
b is 10
.
y-intercept is at (0,10)